### v1

# Finite-temperature quantum topological order in three dimensions

Abstract
We identify a three-dimensional system that exhibits long-range entanglement at sufficiently small but nonzero temperature—it therefore constitutes a quantum topological order at finite temperature. The model of interest is known as the fermionic toric code, a variant of the usual 3D toric code, which admits emergent fermionic point-like excitations. The fermionic toric code, importantly, possesses an anomalous 2-form symmetry, associated with the space-like Wilson loops of the fermionic excitations. We argue that it is this symmetry that imbues low-temperature thermal states with a novel topological order and long-range entanglement. Based on the current classification of three-dimensional topological orders, we expect that the low-temperature thermal states of the fermionic toric code belong to an equilibrium phase of matter that only exists at nonzero temperatures. We conjecture that further examples of topological orders at nonzero temperatures are given by discrete gauge theories with anomalous 2-form symmetries. Our work therefore opens the door to studying quantum topological order at nonzero temperature in physically realistic dimensions.
In recent decades, one of the central goals of condensed matter physics has been the classification and characterization of quantum phases of matter at zero temperature. This has led, most notably, to the concept of topological order, where long-range entanglement manifests in exotic properties, such as anyonic excitations. The investigation of zero-temperature topological orders has culminated in conjectured classifications in low dimensions based on sophisticated mathematics Kitaev (2006); Wen (2015); Johnson-Freyd (2022); Wen (2017); Lan et al. (2018); Lan and Wen (2019).
Despite the theoretical progress, physical experiments cannot be performed in complete isolation from their environment. Consequently, systems will inevitably equilibrate to thermal states at nonzero temperature—possibly destroying the zero-temperature order  [^1] Topological orders in low dimensions can be stabilized far from thermal equilibrium, but this requires highly engineered dynamics (e.g. active error correction). This may be feasible in quantum devices but is less so in conventional solid state systems.. A question of fundamental importance is thus: what topological phases of matter can be realized by thermal states above zero temperature?
It is known that certain topological orders can survive at nonzero temperature; however, to date, all known examples require four spatial dimensions or greater Dennis et al. (2002); Alicki et al. (2010); Hastings (2011); Bombin et al. (2013); Hsin et al. (2024). The prospect of finite-temperature topological order in fewer dimensions has been bleak, with no-go theorems in Refs. Nussinov and Ortiz (2008); Bravyi and Terhal (2009); Alicki et al. (2009); Yoshida (2011); Roberts et al. (2017); Kato and Brandão (2019); Kochanowski et al. (2024); Bakshi et al. (2025); Rouzé et al. (2024). This is usually attributed to the fact that these phases possess point-like excitations, which proliferate above zero temperature and destroy the topological order Peierls (1936); Lan et al. (2018); Lan and Wen (2019); Castelnovo and Chamon (2008).
Three-dimensional topological orders, such as the 3D toric code, however, exhibit both point-like excitations and loop-like excitations. The latter of which only proliferate at some finite critical temperature Peierls (1936); Hastings et al. (2014), implying a finite-temperature phase transition. It is tempting to conclude that the topological order persists at nonzero temperature Castelnovo and Chamon (2008). However, this is not guaranteed to be quantum topological order at finite temperature, which is characterized by long-range entangled (LRE) thermal states, according to the definitions in Refs. Hastings (2011); Sang et al. (2024).
Indeed, for the 3D toric code, it was shown in Ref. Hastings (2011) that its finite-temperature thermal states are short-range entangled (SRE), i.e., they are classical mixtures of SRE pure states. Therefore, its finite-temperature phase transition is between classical phases Lu et al. (2020). It was suggested in Ref. Hastings (2011) that the same conclusion holds for all finite group gauge theories in 3D. Given that these exhaust a large class of topological orders in 3D Lan et al. (2018); Lan and Wen (2019), it is important to settle: is it possible to have quantum topological order at nonzero temperature in 3D?
In this letter, we answer this question in the affirmative. We identify a Hamiltonian in 3D with local interactions such that, below a finite-temperature, its thermal states are topologically ordered, and moreover cannot be approximated by SRE states. It therefore constitutes the first example of a finite-temperature topological order in fewer than four spatial dimensions.
Our example is based on the so-called fermionic toric code (f TC), whose zero-temperature topological order is explored in Refs. Levin and Wen (2003); von Keyserlingk and Burnell (2015); Chen and Kapustin (2019); Chen et al. (2021). Similar to the 3D toric code, referred to henceforth as the bosonic toric code (bTC), the f TC is defined on a system of qubits. The key difference from the bTC is that its elementary point-like excitations have fermionic exchange statistics – i.e., swapping the positions of two of the excitations produces a sign in the pure state wave function. The model furthermore possesses an anomalous 2-form symmetry—simply saying that it is invariant under moving a fermionic excitation along a closed path Bhardwaj et al. (2017); Chen et al. (2018); Chen and Kapustin (2019); Chen et al. (2021); Ellison et al. (2023). The symmetry is considered to be ‘anomalous’, because the resulting string operator inherits the fermionic self-statistics. We show that this is directly responsible for the long-range entanglement of the low-temperature thermal states.
We emphasize that, although the f TC can be used as a quantum memory at zero temperature, it strictly behaves as a classical memory for nonzero temperatures below a finite-temperature phase transition. This is in contrast to the 4D toric code, which serves as a quantum memory at low temperatures Alicki et al. (2010).
We begin with a description of the f TC before proving that its low-temperature thermal states are LRE, as made precise in Theorem 1.
Figure 1: Graphical representations of (a) a vertex term $A_{v}$, (b) plaquette terms $B_{p}$, and (c) the fermion short string operators $W_{e}$. Pauli $X$ and $Z$ operators are represented by red and blue edges, respectively.
Fermionic toric code – The f TC is defined on a cubic lattice with a single qubit on each edge, and for simplicity, we assume that the system has periodic boundary conditions. The Hamiltonian can be written as:
$H=-\sum_{v}A_{v}-\sum_{p}B_{p},$ (2)
where the first sum is over vertices $v$ and the second is over plaquettes $p$. We refer to the $A_{v}$ and $B_{p}$ as the vertex and plaquette terms, respectively. They are represented graphically in Fig. 1.
The Hamiltonian terms are mutually commuting and thus define a Pauli stabilizer code. They further satisfy the following relations on a torus:
$\prod_{v}A_{v}=1,\quad\prod_{p\in c}B_{p}=1,$ (3)
where the second product is over the six plaquettes of an arbitrary cube $c$. These relations are identical to those of the bTC, and as such, the models share the same finite-temperature phase transition Weinstein et al. (2020).
Excitations– The f TC exhibits two types of excitations, which correspond to violations of the vertex and plaquette terms. The violations of the $A_{v}$ terms are point-like and live at the vertices of the lattice. These excitations can be created by operators supported along a path in the lattice, i.e., they are created at the endpoints of string operators. The string operators can be built from the short string operators $W_{e}$ shown in Fig. 1. Notably, in contrast to the bTC, the point-like excitations of the f TC have fermionic exchange statistics, as shown in Ref. Chen and Kapustin (2019); Chen et al. (2021).
It is important to note that the Hamiltonian commutes with the string operators defined on closed paths. This defines a 2-form symmetry of the Hamiltonian, i.e., where the symmetry operators are supported on codimension-2 submanifolds. Due to the point-like excitations having fermionic exchange statistics, the 2-form symmetry is anomalous. In other words, there is an obstruction to gauging the symmetry, as this would correspond to condensing the emergent fermions. Note, however, that the symmetry is anomaly free in the presence of physical fermions as these can be bound to the emergent fermions and condensed Chen et al. (2021).
The second type of excitation in the f TC corresponds to violations of the $B_{p}$ terms. Because of the local relations in Eq. (3), every cube must have an even number of violated $B_{p}$ terms. This means that the violated $B_{p}$ terms form loops in the dual lattice. Accordingly, these excitations are loop-like excitations.
The loop excitations can be created at the boundaries of membrane operators supported on surfaces in the dual lattice. One choice of membrane operator for a surface $\Sigma$ takes the form $M_{\Sigma}=\prod_{e\in\Sigma}X_{e}$. This operator commutes with all of the $A_{v}$ terms and only fails to commute with the $B_{p}$ terms at the boundary of $\Sigma$.
In analogy to the bTC, the point-like excitations and the loop-like excitations have nontrivial mutual braiding statistics, i.e., moving a fermion around a loop excitation incurs a sign. In other words, string operators that move a fermion along a closed path detect the number of loop excitations enclosed modulo 2.
LRE at finite temperature – We now establish the main result of this work: the low-temperature thermal states of the f TC are LRE, as stated more precisely below:
Theorem 1.
Below a finite temperature $T_{0}$, the fidelity  [^2] We take the fidelity of two mixed states $\sigma_{1}$ and $\sigma_{2}$ to be: $F(\sigma_{1},\sigma_{2})=\operatorname{Tr}\left[\sqrt{\sqrt{\sigma_{1}}\sigma_{2}\sqrt{\sigma_{1}}}\right]^{2}.$ between the thermal state $\rho_{\rm{f\hskip 0.35002ptTC}}$ and an arbitrary SRE state $\sigma_{\rm SRE}$ [see Eq. (9)] satisfies
$F(\rho_{\rm{f\hskip 0.35002ptTC}},\sigma_{\rm SRE})=O(L^{-\infty}),$ (4)
i.e., the fidelity decays faster than any polynomial in the system size $L$. This implies, via the Fuchs-van de Graaf inequalities, that the trace distance satisfies:
$\|\rho_{\rm{f\hskip 0.35002ptTC}}-\sigma_{\rm SRE}\|_{1}=1-O(L^{-\infty}).$ (5)
We start with an outline for the proof of Theorem 1. The temperature $T_{0}$ is chosen to be low enough that the large loop excitations are suppressed in $\rho_{\text{f\nobreak\hskip 0.56001ptTC}}$. As such, the majority of the loop excitations can be removed with a quasi-local channel (QLC)  [^3] We take a QLC to be a quantum channel whose dilation is a quasi-local unitary. By a quasi-local unitary, we mean a quantum circuit for which the depth and the size of the supports of the gates are at most polylogarithmic in the system size.. This yields a mixed state that well approximates a state with exactly no loop excitations, denoted as $\rho_{\varnothing}$.
Next, we leverage the fact that the loop-less state $\rho_{\varnothing}$ possesses the anomalous 2-form symmetry of the f TC as a strong symmetry. Meaning that, for any symmetry operator $W$ supported on a contractible path, $\rho_{\varnothing}$ satisfies $W\rho_{\varnothing}=\rho_{\varnothing}$. This is because the symmetry merely detects the parity of loop excitations enclosed by the path, of which there are none for $\rho_{\varnothing}$. The anomaly then forbids there from being any SRE state capable of approximating $\rho_{\varnothing}$, as shown recently in Ref. Li et al. (2024).
Finally, we use that the QLC that cleans up the loop excitations of $\rho_{{\text{f\nobreak\hskip 0.56001ptTC}}}$ maps SRE states to SRE states. If $\rho_{{\text{f\nobreak\hskip 0.56001ptTC}}}$ were well approximated by a SRE state, we would be able to clean up the loop excitations to obtain a SRE approximation to $\rho_{\varnothing}$ – a contradiction. In the remainder of this letter, we clarify and fill in the details of the proof. The argument proceeds in three steps.
Step one: The first step is to define a QLC that approximately maps the thermal state $\rho_{\text{f\nobreak\hskip 0.56001ptTC}}$ to the loop-less state $\rho_{\varnothing}$. We accomplish this with a QLC $\mathcal{C}$, composed of five layers of QLCs. In each layer, we clean up loop excitations on disjoint patches of linear size at least $\ln(L)^{2}$. After applying all five layers, the only loop excitations that are left behind must have lengths larger than $\ln(L)^{2}$. These large loop excitations are exponentially suppressed in the thermal state at low temperatures. Thus, the QLC approximately maps $\rho_{\text{f\nobreak\hskip 0.56001ptTC}}$ to $\rho_{\varnothing}$, in terms of the fidelity of mixed states. We note that, alternatively, a QLC based on the sweep decoder of Ref. Kubica and Preskill (2019) can be used to remove the small loop excitations, as elaborated on in the Supplemental Materials (SM).
Figure 2: 2D representation of the QLC $\mathcal{C}$. (a) A configuration of loop excitations, shown in red. (b) To define the layer $\mathcal{C}_{1}$ of the QLC, we divide the system into blocks of linear size at least $\ln(L)^{2}$ (dashed lines). The loop excitations that are contained entirely within a block are removed with a unitary $U_{\mathcal{L}_{b}}$, as in Eq. (6). These loop excitations are shown in blue. (c) The layer $\mathcal{C}_{2}$ is defined by blocking the system (dashed lines) into blocks centered on the vertical seams of the initial blocking (solid lines). The vertices of the initial blocking are marked with a black dot. The loop excitations shown in blue are entirely contained within the second blocking and are removed by $\mathcal{C}_{2}$. (d-e) Similarly, the third and fourth layers $\mathcal{C}_{3},\mathcal{C}_{4}$ are specified by blocking the system into blocks centered on the horizontal seams or vertices, respectively, of the initial blocking. The remaining loops (red) must have length greater than $\ln(L)^{2}$.
To define $\mathcal{C}$, we begin by partitioning the system into disjoint blocks of linear size $\ln(L)^{2}$, shown for a 2D system in Fig. 2(b). On each of the blocks we apply a quantum channel that removes all loop excitations that are entirely contained within the block. The quantum channel $\mathcal{C}_{b}$ on a block $b$ acts as:
$\mathcal{C}_{b}(\cdot)=\sum_{\mathcal{L}|_{b}}U_{\mathcal{L}_{b}}P_{\mathcal{L}|_{b}}\cdot P_{\mathcal{L}|_{b}}U_{\mathcal{L}_{b}},$ (6)
where $\mathcal{L}|_{b}$ is a loop configuration restricted to $b$, $P_{\mathcal{L}|_{b}}$ is the projector onto the configuration $\mathcal{L}|_{b}$, and $U_{\mathcal{L}_{b}}$ is a product of $X$ operators that cleans up all of the loop excitations $\mathcal{L}_{b}$ that are fully contained within $b$. The first layer of the QLC $\mathcal{C}$ is thus $\mathcal{C}_{1}=\bigotimes_{b}\mathcal{C}_{b}$.
Notice that loop excitations that cross between blocks are not removed by $\mathcal{C}_{1}$, regardless of their size. Therefore, in the subsequent layers, we clean up loop excitations along the seams of the first blocking, starting with those perpendicular to the $x$-axis, referred to as $x$-seams. We again partition the system into blocks of linear size $\ln(L)^{2}$, but this time, the blocks are chosen so that they are centered on the $x$-seams of the first blocking, as depicted in Fig. 2(c). We take the second layer of $\mathcal{C}$ to be analogous to $\mathcal{C}_{1}$ but for the new choice of blocks. We refer to this QLC as $\mathcal{C}_{2}$. We repeat this process for the $y$-seams, $z$-seams, and for the corners of the initial blocking. This defines the QLCs $\mathcal{C}_{3}$, $\mathcal{C}_{4}$, and $\mathcal{C}_{5}$.
Altogether, the QLC $\mathcal{C}$ is ${\mathcal{C}_{5}\circ\mathcal{C}_{4}\circ\mathcal{C}_{3}\circ\mathcal{C}_{2}\circ\mathcal{C}_{1}}$. This channel is guaranteed to clean up all of the loop excitations with length less than $\ln(L)^{2}$. In the SM, we bound the fidelity of $\mathcal{C}(\rho_{\text{f\nobreak\hskip 0.56001ptTC}})$ and the state $\rho_{\varnothing}$ as
$F(\mathcal{C}(\rho_{{\text{f\nobreak\hskip 0.56001ptTC}}}),\rho_{\varnothing})\geq 1-L^{3}e^{-(2\beta-\ln(5))\ln(L)^{2}}.$ (7)
Therefore, if the inverse temperature $\beta$ satisfies $\beta>\frac{\ln(5)}{2}$, the fidelity approaches unity faster than any polynomial in $L$. Below a temperature $T_{0}=2\ln(5)^{-1}$, we have:
$F(\mathcal{C}(\rho_{{\text{f\nobreak\hskip 0.56001ptTC}}}),\rho_{\varnothing})=1-O(L^{-\infty}).$ (8)
Note that the temperature $T_{0}\approx 1.24$ is below the critical temperature $T_{c}\approx 1.31$ found in Ref. Castelnovo and Chamon (2008).
Step two: The second step is to argue that $\rho_{\varnothing}$ cannot be approximated by a SRE state. To make this explicit, we let $\sigma_{\text{SRE}}$ be a mixture of SRE states, such as:
$\sigma_{\text{SRE}}=\sum_{i}p_{i}|\psi_{\text{SRE}}^{i}\rangle\langle\psi_{\text{SRE}}^{i}|.$ (9)
We assume here that each $|\psi_{\text{SRE}}^{i}\rangle$ can be prepared from a product state with a local unitary circuit with depth less than $d$, where $d$ is polylogarithmic in the system size. Our goal is to then bound the fidelity between $\rho_{\varnothing}$ and $\sigma_{\text{SRE}}$ as a function of the depth $d$.
The fidelity can be immediately bounded using a simple generalization of Corollary 2 in Ref. Li et al. (2024) to 3D. Corollary 2 gives us the bound:
$F(\rho_{\varnothing},\sigma_{\text{SRE}})<e^{\frac{-cL^{3}}{(d+1)^{3}}},$ (10)
for some positive constant $c$. Note that the only difference from the explicit bounds in Ref. Li et al. (2024) are the powers of 3, which are a consequence of 3D. If $d$ is polylogarithmic in $L$, we have our desired result:
$F(\rho_{\varnothing},\sigma_{\text{SRE}})=O(L^{-\infty}).$ (11)
The absence of loop excitations in $\rho_{\varnothing}$ means that it obeys a strong 2-form symmetry. Thus, while the thermal state $\rho_{\text{f\nobreak\hskip 0.56001ptTC}}$ itself is not exactly symmetric, by Eq. (11), the symmetry can be restored (approximately) by cleaning up the loop excitations. Importantly, this symmetry is anomalous, due to the self-statistics of the fermions, which allows us to employ the results of Ref. Li et al. (2024). This is notably distinct from the bTC, whose 2-form symmetry is anomaly free. We again emphasize that in the presence of physical fermions, the 2-form symmetry of the f TC also becomes anomaly free, and hence, $\rho_{\varnothing}$ can be written as a mixture of SRE states with physical fermions.
Step three: We now upper bound the fidelity between a thermal state $\rho_{\text{f\nobreak\hskip 0.56001ptTC}}$ below the temperature $T_{0}$ and an arbitrary SRE state $\sigma_{\text{SRE}}$. To begin, we notice that the quantum channel $\mathcal{C}$ that cleans up small loop excitations can only increase the fidelity, according to the data processing inequality, i.e.,
$F(\rho_{\text{f\nobreak\hskip 0.56001ptTC}},\sigma_{\text{SRE}})\leq F\left(\mathcal{C}(\rho_{\text{f\nobreak\hskip 0.56001ptTC}}),\mathcal{C}(\sigma_{\text{SRE}})\right).$ (12)
Since $\mathcal{C}$ is a QLC, the state $\mathcal{C}(\sigma_{\text{SRE}})$ is also SRE, up to a tensor product with the identity, as spelled out in the SM. In particular, there exists a SRE state of the form $\tilde{\sigma}_{\text{SRE}}=\mathcal{C}(\sigma_{\text{SRE}})\otimes I$. Taking a tensor product with the identity does not change the fidelity, so we may write:
$F(\rho_{\text{f\nobreak\hskip 0.56001ptTC}},\sigma_{\text{SRE}})\leq F\left(\mathcal{C}(\rho_{\text{f\nobreak\hskip 0.56001ptTC}}),\tilde{\sigma}_{\text{SRE}}\right).$ (13)
The bounds in Eqs. (8) and (11) suggest that $\mathcal{C}(\rho_{\text{f\nobreak\hskip 0.56001ptTC}})$ is close to $\rho_{\varnothing}$ and that $\rho_{\varnothing}$ is far from any SRE state. This then implies that $\mathcal{C}(\rho_{\text{f\nobreak\hskip 0.56001ptTC}})$ must be far from any SRE state. This intuition can be made precise using the triangle inequality for the trace distance and the Fuchs-van de Graaf inequalities Fuchs and van de Graaf (1999). Combining these two, arbitrary mixed states $\sigma_{1}$, $\sigma_{2}$ and $\sigma_{3}$ must satisfy:
$F(\sigma_{2},\sigma_{3})\leq 1-\left[1-\sqrt{F(\sigma_{1},\sigma_{3})}-\sqrt{1-F(\sigma_{1},\sigma_{2}))}\right]^{2}.$
Setting $\sigma_{1}=\rho_{\varnothing}$, $\sigma_{2}=\mathcal{C}(\rho_{\text{f\nobreak\hskip 0.56001ptTC}})$, and $\sigma_{3}=\tilde{\sigma}_{\text{SRE}}$ and plugging in the inequalities of Eqs. (8) and (11), we obtain
$F(\mathcal{C}(\rho_{\text{f\nobreak\hskip 0.56001ptTC}}),\tilde{\sigma}_{\text{SRE}})=O(L^{-\infty}).$ (14)
Together with Eq. (13), this gives us our main result and concludes the proof of Theorem 1:
$F(\rho_{\text{f\nobreak\hskip 0.56001ptTC}},\sigma_{\text{SRE}})=O(L^{-\infty}).$ (15)
Conclusions – We have demonstrated the existence of finite-temperature quantum topological order in three dimensions, exemplified by the f TC below a temperature ${T_{0}=2\ln(5)^{-1}}$. The crucial property of the f TC, which enforces long-range entanglement at nonzero temperature, is that it posses an anomalous 2-form symmetry. This is in sharp contrast to the bTC, which has been the focus of previous investigations into finite-temperature topological order in three dimensions Castelnovo and Chamon (2008); Hastings (2011); Lu et al. (2020).
Our argument used that $\rho_{\text{f\nobreak\hskip 0.56001ptTC}}$ at low temperatures can be related to the loop-less state $\rho_{\varnothing}$ by a QLC. We conjecture that the opposite is also true, and there exists a QLC mapping $\rho_{\varnothing}$ back to $\rho_{\text{f\nobreak\hskip 0.56001ptTC}}$. This would imply that the two belong to the same mixed-state phase of matter, by the definitions of Refs. Sang et al. (2024); Coser and Pérez-García (2019). Furthermore, due to arguments similar to Refs. Wang et al. (2025); Sohal and Prem (2025); Ellison and Cheng (2025), we expect $\rho_{\varnothing}$ to be inequivalent to any pure state topological order. This would then suggest that the finite-temperature f TC order that we have identified is a quantum phase of matter that exists only at nonzero temperature, or in open systems, as a mixed-state topological order Fan et al. (2024); Bao et al. (2023); Wang et al. (2025); Sohal and Prem (2025); Ellison and Cheng (2025); Coser and Pérez-García (2019); Li and Mong (2024); Sang and Hsieh (2024); Zhang et al. (2025); Ogata (2025).
We expect that further examples of finite-temperature topological orders can be found by generalizing the f TC to other models with anomalous 2-form symmetries. A large class of examples are given by Walker-Wang models based on premodular categories with a transparent fermion Walker and Wang (2011); Deligne (2002); Ellison and Cheng (2025). At an intuitive level, these are discrete $G$ gauge theories obtained by gauging the symmetry of a system with physical fermions.
The existence of finite-temperature topological order in 3D opens up a number of interesting questions. First, is the long-range entanglement exhibited by the f TC thermal states, and thus the topological order, stable to perturbations of the Hamiltonian? Our arguments used the fact that the f TC Hamiltonian is 2-form symmetric. For a perturbed Hamiltonian, however, the 2-form symmetry is in general emergent, i.e., the system is only guaranteed to have an anomalous 2-form symmetry in its ground-state subspace. Can QLCs be developed to clean up the loop excitations in this case, to (approximately) restore the anomalous 2-form symmetry of the ground state?
Another pressing question is: what diagnostics can be used to detect the topological order of the f TC at nonzero temperature? In principle, one could use a QLC to clean up the loop excitations, then measure the anomalous 2-form symmetry operators in the loop-less state. Alternatively, it is natural to expect that the long-range entanglement can be detected by computing the topological entanglement negativity, analogous to the 4D toric code in Ref. Lu et al. (2020). We leave the search for efficiently computable diagnostics of the finite-temperature topological order to future explorations.
Furthermore, it would be interesting to identify physical systems able to realize the finite-temperature topological order of the f TC. The Kitaev model in Ref. Mandal and Surendran (2009) may be particularly amenable to this pursuit, due to it natively requiring only 2-body interactions. Similar Kitaev-like 2-body interactions have been proposed for neutral atom systems through Floquet engineering in Ref. Kalinowski et al. (2023).
Acknowledgments – TDE is grateful to Timothy H. Hsieh, Shengqi Sang, and Jinmin Yi for valuable conversations about SRE states. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. CvK is supported by UKRI FLF grant MR/Z000297/1. MC is supported by NSF grant DMR-2424315. TR was supported by Stanford’s Q-FARM Bloch Postdoctoral Fellowship in Quantum Science and Engineering and by the HUN-REN Hungarian Research Network through the HUN-REN Welcome Home and Foreign Researcher Recruitment Programme 2023 and through the Supported Research Groups Programme, HUN-REN-BME-BCE Quantum Technology Research Group (TKCS-2024/34).